Find the centroid of the triangle whose vertices are (−1,0), (1,0), 𝑎𝑛𝑑 (0,3


1
Expert's answer
2022-05-15T10:48:59-0400

We know that the centroid of the triangle "(\frac{x_1+x_2+x_3}{3};\frac{y_1+y_2+y_3}{3})"

Where "(x_1;y_1);(x_2;y_2);(x_3;y_3)" are the vertices of the

triangle SO, centroid= "(\frac{(-1)+1+0}{3};\frac{0+0+3}{3})=(\frac{0}{3};\frac{3}{3})=(0;1)"

Hence, this is the answer.